ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 43A62, 43A46 Schwartz, Alan, University of Missouri-St. Louis, USA, Gebuhrer, Olivier, Universite Louis Pasteur, Strasbourg, France Hypergroups and their maximum subgroup A hypergroup is a probability preserving measure algebra with certain additional properties so that it resembles the measure algebra of a topological group; examples include the algebra of bi-invariant measures on a group; other examples arise in connection with product formulas for special functions. Every hypergroup contains a maximal topological subgroup in a natural way. The purpose of this article is to show how homomorphisms, idempotents, and Sidon sets of a compact commutative hypergroup can often be described in terms of those of the maximum subgroup. Examples are provided.